OpenStudy (anonymous):

decide whether the equation is a trigonometric identity. sin^2(x)csc^2(x)=sin^2(x)+cos^2(x)

OpenStudy (zzr0ck3r):

What do we know that \(\cos^2(x)+\sin^2(x)=?\)

OpenStudy (zzr0ck3r):

@Reid448

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (zzr0ck3r):

and \[\sin^2(x)*\csc^2(x)= \sin^2(x) * \frac{1}{\sin^2(x)}=?\]

OpenStudy (anonymous):

and \[\sin ^{2}=\frac{ 1-\cos2x }{ 2 }\]

OpenStudy (anonymous):

what you said also equals 1 so that makes this an identity

OpenStudy (anonymous):

thanks very much

OpenStudy (dinamix):

@zzr0ck3r i want ask u question how proof it sin^2(x) + cos^2(x) to---> sin^2(x) csc^2(x) cuz its relation equivalent

OpenStudy (zzr0ck3r):

?

OpenStudy (dinamix):

u understand my question or no ?

OpenStudy (zzr0ck3r):

I don't think so.

OpenStudy (dinamix):

i mean u proof it from one side only

OpenStudy (zzr0ck3r):

we showed both sides equal 1

OpenStudy (zzr0ck3r):

you understand @dinamix